Consider the nonlinear system
?x′ = ln(y^2 − x) and y'=x-y-1
(a)Find all the critical points

(b)Find the corresponding linearized system near the critical points.

(c) Classify the (i) type (node, saddle point, · · · ), and (ii) stability of the critical points for the corresponding linearized system.

(d) What conclusion can you obtain for the type and stability of the critical points for the original nonlinear system?
Answer #1

ebolution Given that , x- lalya-x) = f(x,y) y= n-y-1 glug) (a) f(x, y) = 0 = y2-1=1 => x=42-1 gun,y)= 0 = x-y-1 =) 4²-4-2=0= ( b ) fy y?re -1 fore y²-a Ja = 1 gy four (0,-). Coorousbonding linear System us fa 10,-1). U + fy (0, -1). V V = 9,10,-1)fou 13.21, A: E - 1A-2 I)= 0 = (-1-752- 4= 0 22 +29-3=0. = (2+3) (2-1) = 0 n=1 1-3 Obbosite sign. Saddle boint always unstabl